OpenStudy (anonymous):

Phillip deposited $5,670 into a savings account 19 years ago. The account has an interest rate of 4.1% and the balance is currently $12,356.04. How often does the interest compound? daily monthly quarterly annually Please help me. I can't fail this test.

jimthompson5910 (jim_thompson5910):

Hint: A = P(1+r/n)^(n*t) 12356.04 = 5670(1+0.041/n)^(n*19) I'll let you finish and solve for n

OpenStudy (anonymous):

I did do that, exactly that. However, it never equaled 12356.04, so that's why I got confused. My brain is fried so let me try again, I probably made a mistake. Hold on...

jimthompson5910 (jim_thompson5910):

alright

jimthompson5910 (jim_thompson5910):

What you could do instead is plug in each possible answer choice and see which ones are true or false

jimthompson5910 (jim_thompson5910):

but I'll see what you get here

OpenStudy (anonymous):

Wow, how stupid. The answer is daily :)

jimthompson5910 (jim_thompson5910):

yep because plugging in n = 365 gives you roughly 12,356.0444063578 which rounds to 12,356.04

jimthompson5910 (jim_thompson5910):

plugging in n = 4 for example will give you 12,307.6844083382, which is slightly off

OpenStudy (anonymous):

Yep, when I typed that in my calculator, it gave me a bizarre answer. When I used an online calculator, it gave me 12,356.04. Thank you SO much for your help. Can you help me with ONE more please? :)

jimthompson5910 (jim_thompson5910):

sure thing, glad to be of help and go for it

jimthompson5910 (jim_thompson5910):

what's your other question?

OpenStudy (anonymous):

Thank you so much!! A video game has been developed whereby the player must earn a certain number of experience points before progressing to the next level. On level 1, a player must earn 1 experience point. On level 2, a player must earn 6 experience points. On level 3, a player must earn 13 experience points. On level 4, a player must earn 22 experience points. How many experience points must a player earn on level 7? 33 46 54 61

OpenStudy (anonymous):

Every question that is like this I get wrong. I just don't understand it at all.

jimthompson5910 (jim_thompson5910):

this one is a bit tricky, but... the jump from 1 to 6 is 5 (since 1+5 = 6) the jump from 6 to 13 is 7 (since 6+7 = 13) the jump from 13 to 22 is 9 (since 13+9 = 22) so we start off by adding 5, then we add 7, then we add 9, etc The pattern is +5, +7, +9, ...

jimthompson5910 (jim_thompson5910):

basically we add on odd numbers from 5 onward

jimthompson5910 (jim_thompson5910):

The next odd number after 9 is 11, so we add 11 to the last term of 22 to get 22+11 = 33 This is the 5th term

jimthompson5910 (jim_thompson5910):

The 6th term is found by adding 13 (the next odd after 11) to 33 to get 33+13 = 46 Finally, the 7th term is 46+15 = 61

jimthompson5910 (jim_thompson5910):

So you need 61 points on the 7th level

OpenStudy (anonymous):

So 61 would be the 7th term?

jimthompson5910 (jim_thompson5910):

The good news is that you can interpolate a polynomial (a quadratic to be specific) to make things a bit shorter, but luckily the 7th term isn't too far away

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

is this site glitching on you a bit?

jimthompson5910 (jim_thompson5910):

it says I'm typing when I'm not lol

OpenStudy (anonymous):

Thank you so much! I think I finally understand it. You are such a great help, I can't thank you enough!